OpenStudy (anonymous):

3√343+3/4 3√-8 A) -5 1/2 B) 1/2 C) 8 1/2 need help.

OpenStudy (triciaal):

should that be the cube root of 343 plus 3/4 cube root of -8?

OpenStudy (triciaal):

cube root of 343 = 7 and cube rt of -8 = -2 7 + 3/4(-2) = 7 - 3/2 = 5 1/2 was this choice D

OpenStudy (triciaal):

are you there?